# Itô’s formula for flows of conditional measures on semimartingales

### v2

@@ 2.1. Assumptions
For $p\geq 2$, a function $\Phi(\mu,\bm{y})$ is a $C^{2,2}({\mathcal{P}}_{p}({\mathbb{R}}^{d})\times{\mathbb{R}}^{d})$ if for any $\mu\in\mathcal{P}_{p}({\mathbb{R}}^{d})$ and $\bm{y}\in{\mathbb{R}}^{d}$, there exists a continuous mapping $(\mu,\bm{y},\bm{x}_{1})\to\frac{\delta\Phi}{\delta\mu}(\mu,\bm{y},\bm{x}_{1})$, and there exists a continuous mapping $(\mu,\bm{y},\bm{x}_{1},\bm{x}_{2})\to\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}(\mu,\bm{y},\bm{x}_{1},\bm{x}_{2})$ that is symmetric in its two arguments and with the following properties,
  • continuously differentiable: $\nabla_{\bm{y}}\Phi(\mu,\bm{y})$, $\nabla_{\bm{y}}^{2}\Phi(\mu,\bm{y})$, $\frac{\delta\Phi}{\delta\mu}(\mu,\bm{y},\bm{x}_{1})$, $\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}(\mu,\bm{y},\bm{x}_{1})$, $\nabla_{\bm{y}}\frac{\delta\Phi}{\delta\mu}(\mu,\bm{y},\bm{x}_{1})$ $\nabla^{2}_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}(\mu,\bm{y},\bm{x}_{1})$, $\nabla_{\bm{x}_{1}}\nabla_{\bm{y}}\frac{\delta\Phi}{\delta\mu}(\mu,\bm{y},\bm{x}_{1})$, $\nabla_{\bm{x}_{1}}\nabla_{\bm{x}_{2}}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}(\mu,\bm{y},\bm{x}_{1},\bm{x}_{2})$ all exist and are continuous functions for all $\bm{y},\bm{x},\bm{x}^{\prime}\in{\mathbb{R}}^{d}$, $\mu\in{\mathcal{P}}_{p}({\mathbb{R}}^{d})$,
@@
  • uniform polynomial-growth : there exists a constant $c>0$ such that for all $\bm{x}_{1},\bm{x}_{2},\bm{y}\in{\mathbb{R}}^{d}$, $\mu\in{\mathcal{P}}({\mathbb{R}}^{d})$, we have
@@
    $\bigg{|}\nabla^{2}_{\bm{x}}\frac{\delta\Phi}{\delta\mu}\bigg{|}+\bigg{|}\nabla_{\bm{x}_{1}\bm{y}}\frac{\delta\Phi}{\delta\mu}\bigg{|}$ $\leq c\big{(}1+|\bm{x}_{1}|^{p-2}+|\bm{y}|^{{p-2}}\big{)},$
    $\bigg{|}\nabla_{\bm{x}_{1},\bm{x}_{2}}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\bigg{|}$ $\leq c\big{(}1+|\bm{x}_{1}|^{p-2}+|\bm{x}_{2}|^{p-2}+|\bm{y}|^{p-2}\big{)},$
  • fundamental theorem of calculus: for everything $\mu,\nu\in{\mathcal{P}}_{2}({\mathbb{R}}^{d})$, we have
@@
    $\Phi(\bm{x}_{0},\mu)-$ $\Phi(\bm{x}_{0},\nu)=\int_{0}^{1}\int_{{\mathbb{R}}^{d}}\frac{\delta\Phi}{\delta\mu}\big{(}\nu,\bm{y},\bm{x}_{1}\big{)}(\mu-\nu)({\mathrm{d}}\bm{x}_{1}){\mathrm{d}}\lambda$
    $+\int_{0}^{1}\int_{0}^{t}\int_{{\mathbb{R}}^{d}}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}s\mu+(1-s)\nu,\bm{y},\bm{x}_{1},\bm{x}_{2}\big{)}(\mu-\nu)({\mathrm{d}}\bm{x}_{1})(\mu-\nu)({\mathrm{d}}\bm{x}_{2}){\mathrm{d}}s{\mathrm{d}}t.$
Such $\frac{\delta\Phi}{\delta\mu}$ and $\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}$ are called (a version of) the linear derivative and the second derivative of $\Phi$.
@@ 2.2. Main result
Given a completed filtered probability space $(\Omega,{\mathcal{F}},{\mathbb{F}}=({\mathcal{F}}_{t})_{t\in[0,T]},{\mathbb{P}})$, which supports two ${\mathbb{R}}^{d}$-valued ${\mathbb{F}}$-semimartingales $\bm{X}=(\bm{X}_{t})_{t\in[0,T]}$ and $\bm{Y}=(\bm{Y}_{t})_{t\in[0,T]}$ satisfying Assumption 2.1. Let $\mu_{t}=$Law$(X_{t}|{\mathcal{G}}_{t})$, with the subfiltration ${\mathbb{G}}=({\mathcal{G}}_{t})_{t\in[0,T]}\subset{\mathbb{F}}$ satisfying Assumption 2.5. Then, for $p\geq 2$ and any $\Phi\in C^{2,2}(\mathcal{P}_{p}({\mathbb{R}}^{d})\times{\mathbb{R}}^{d})$, we have
@@
$\qquad\quad+\frac{1}{2}\nabla^{2}_{\bm{x}_{1}\bm{x}_{2}}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-},\bm{X}^{\prime\prime}_{s-}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime\prime}]^{c}_{t}+\nabla^{2}_{\bm{x}_{1}\bm{y}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{Y}]^{c}_{s}\bigg{)}$
@@
for all $t\in[0,T]$, where $\bm{X}^{\prime},\bm{X}^{\prime\prime}$ are the conditional independent copies of $\bm{X}$ given the sub $\sigma$-algebra ${\mathcal{G}}_{T}$ defined in an enlarged probability space $(\overline{\Omega},\overline{{\mathcal{F}}},\overline{{\mathbb{P}}})$ and $\overline{{\mathbb{E}}}[\cdot|{\mathcal{F}}]$ is the conditional expectation given the extension ${\mathcal{F}}\subset\overline{{\mathcal{F}}}$ in the enlarged probability space.
@@
To ensure that the enlarged space $(\overline{\Omega},\overline{{\mathcal{F}}},\overline{{\mathbb{F}}},\overline{{\mathbb{P}}})$ is properly defined such that $\mu_{t},\bm{X}_{t},\bm{Y}_{t}$ and ${\mathcal{F}}$ are naturally extended in this enlarged space, one needs the notion of independent copies of stochastic processes, as will be developed in the next section and discussed after Corollary 3.2.
3. Conditional independent copy
3.1. Why conditional independent copy?
One of the key components in Theorem 2.7 is the concept of the conditional independent copy of stochastic processes. Here in this section, we illustrate through several examples its properties in the context of flow of conditional laws for semimartingales and its role in the derivation of Itô’s formula.
Let us first recall a mean-field game with common noise where $X_{t}$ is of the form
@@ 3.1. Why conditional independent copy?
where $W,W^{0}$ are two independent Brownian motions and ${\mathbb{G}}=\{{\mathcal{G}}_{t}\}_{t\in[0,T]}$ is the sub filtration generated by $W^{0}$ and ${\mathcal{G}}={\mathcal{G}}_{T}=\sigma(W^{0}_{u},u\leq T)$. By the classical Itô’s formula, the dynamic of $g(X_{t})$ is given by
@@
The above equation follows essentially from the fact that
${\mathbb{E}}\bigg{[}\int_{0}^{t}\eta_{s}{\mathrm{d}}W_{s}\Big{|}{\mathcal{G}}\bigg{]}=0,\qquad{\mathbb{E}}\bigg{[}\int_{0}^{t}\eta_{s}{\mathrm{d}}W^{0}_{s}\Big{|}{\mathcal{G}}\bigg{]}=\int_{0}^{t}{\mathbb{E}}[\eta_{s}|{\mathcal{G}}]{\mathrm{d}}W^{0}_{s},\text{ for all }t\in[0,T]$ (3.2)
see [34, Lemma B.1] for details. Indeed, from (3.1) one can calculate the quadratic variation of $Z_{t}$ to derive the Itô’s formula for $f(Z_{t})$:
@@
where $\widetilde{\sigma}_{t}^{0},\widetilde{X}_{t}$ are conditional respective independent copy of $\sigma^{0}_{t}$ and $X_{t}$ given ${\mathcal{G}}$, which is formally introduced in [11]. Under these expressions, one can derive Itô’s formula for $f(Z_{t})$, which depends on the precise expression of
$\int_{0}^{t}f^{\prime}(Z_{s}){\mathrm{d}}Z_{s}=$ $\int_{0}^{t}f^{\prime}(Y_{s})\bigg{(}{\mathbb{E}}\Big{[}b_{s}g^{\prime}(X_{s})+\frac{1}{2}(\sigma_{s}^{2}+(\sigma^{0}_{s})^{2})g^{\prime\prime}(X_{s})\Big{|}{\mathcal{G}}\Big{]}{\mathrm{d}}s+{\mathbb{E}}\big{[}\sigma^{0}_{t}g^{\prime}(X_{s})|{\mathcal{G}}\big{]}{\mathrm{d}}W^{0}_{s}\bigg{)}$
@@
Now, the issue arises when one considers a general continuous semimartingale $X$ and a general sub $\sigma$-algebra ${\mathcal{G}}$, where the Itô’s formula takes the form:
@@
In this case, in order to characterize the dynamics of $Z_{t}={\mathbb{E}}[g(X_{t})|{\mathcal{G}}]$, one needs an appropriate expression for ${\mathrm{d}}Z_{t}$, ${\mathrm{d}}\langle Z\rangle_{t}$ and more importantly for
$\int_{0}^{t}f^{\prime}(Z_{s}){\mathrm{d}}Z_{s},\text{ and }\int_{0}^{t}f^{\prime\prime}(Z_{s}){\mathrm{d}}\langle Z\rangle_{s}.$ (3.4)
@@
${\mathbb{E}}\bigg{[}\int_{0}^{t}\eta_{s}{\mathrm{d}}X_{s}\Big{|}{\mathcal{G}}\bigg{]}=\int_{0}^{t}{\mathbb{E}}[\eta_{s}|{\mathcal{G}}]{\mathrm{d}}X^{\mathcal{G}}_{s},$ (3.5)
where $X^{\mathcal{G}}_{t}:={\mathbb{E}}[X_{t}|{\mathcal{G}}]$.
Unfortunately this does not hold in general. For instance, take $X_{t}$ from (3.1) with $b_{t}=0$. By (3.2), the left hand side of (3.5) becomes
@@
and (3.5) holds if and only if $\eta_{s}$ and $\sigma_{s}^{0}$ are conditionally uncorrelated of ${\mathcal{G}}$.
This is exactly why the concept of conditional independent copies of stochastic processes: instead of (3.2), let us consider the expression of the following form
$\int_{0}^{t}f^{\prime}(Z_{s}){\mathrm{d}}Z_{s}$ $=\int_{0}^{t}f^{\prime}(Z_{s}){\mathrm{d}}{\mathbb{E}}\big{[}g(X_{s})|{\mathcal{G}}\big{]}=\int_{0}^{t}f^{\prime}(Z_{s}){\mathrm{d}}{\mathbb{E}}\big{[}g(X^{(1)}_{s})|{\mathcal{F}}\big{]}={\mathbb{E}}\bigg{[}\int_{0}^{t}f^{\prime}(Z_{s}){\mathrm{d}}g(X^{(1)}_{s})\Big{|}{\mathcal{F}}\bigg{]}$
$={\mathbb{E}}\bigg{[}\int_{0}^{t}f^{\prime}(Z_{s})\Big{(}g^{\prime}(X^{(1)}_{s}){\mathrm{d}}X^{(1)}_{s}+\frac{1}{2}g^{\prime\prime}(X^{(1)}_{s}){\mathrm{d}}\langle X^{(1)}\rangle_{s}\Big{)}\Big{|}{\mathcal{F}}\bigg{]},$
@@
$\int_{0}^{t}f^{\prime\prime}(Z_{s}){\mathrm{d}}\langle Z\rangle_{s}$ $=\int_{0}^{t}f^{\prime}(Z_{s}){\mathrm{d}}\langle{\mathbb{E}}\big{[}g(X_{\cdot})|{\mathcal{G}}\big{]}\rangle_{s}={\mathbb{E}}\bigg{[}\int_{0}^{t}f^{\prime}(Y_{s}){\mathrm{d}}\langle g(X^{(1)}_{\cdot}),g(X^{(2)}_{\cdot})\rangle_{s}|{\mathcal{F}}\bigg{]}$
$={\mathbb{E}}\bigg{[}\int_{0}^{t}f^{\prime\prime}(Z_{s})g^{\prime}(X^{(1)}_{s})g^{\prime}(X^{(2)}_{s}){\mathrm{d}}\langle X^{(1)},X^{(2)}\rangle_{s}\Big{|}{\mathcal{F}}\bigg{]},$
where $X^{(1)}$, $X^{(2)}$ are two conditional independent copies of $X$ give ${\mathcal{G}}$. With these expressions, one can analyze (3.4) for a general semimartingale $X$, as will be detailed in (2c) and (2d) of Theorem 3.3.
3.2. Examples
To get some intuition for this notion of conditional independent copies of stochastic processes, let us see some examples when ${\mathcal{F}}$ and the sub $\sigma$-algebra ${\mathcal{G}}$ have some special structure.
Example 1. Let $A,B$ are two independent random variables with ${\mathcal{F}}=\sigma(A,B)$ and ${\mathcal{G}}=\sigma(A)$. Then one can enlarge the probability space to include $A^{(i)}$ such that $A^{(i)}$ are independent and with the same distribution of $A$, and are independent of $B$. Then for any $X=f(A,B)$, the conditional independent copies of $X$ are $X^{(i)}=f(A^{(i)},B)$.
Example 2. Let $X$ be the unique strong solution of the SDE
where $W,W^{(0)}$ are two independent Brownian motions, and $b$, $\sigma$, and $\sigma^{0}$ satisfy appropriate continuity conditions. Let ${\mathcal{F}}_{t}=\sigma\big{(}\{W_{s},W^{(0}_{s}\}_{s\leq t}\big{)}$ and sub $\sigma$-algebra ${\mathcal{G}}=\sigma(W^{(0)})$, then one can enlarge the probability space to include independent Brownian motions $W^{(1)}$ and $X^{(1)}$, with $X^{(1)}$ the unique strong solution of
${\mathrm{d}}X^{(1)}_{t}=b(X^{(1)}_{t}){\mathrm{d}}t+\sigma(X^{(1)}_{t}){\mathrm{d}}W^{(1)}_{t}+\sigma^{(0)}(X^{(1)}_{t}){\mathrm{d}}W^{(0)}_{t},$
and $X^{(1)}$ is the conditional independent copies of $X$ given ${\mathcal{G}}$.
There are more examples of conditional independent copies of stochastic processes in the literature of mean-field games. For instance, when defining a semimartingale by stochastic differential equations involving both idiosyncratic and common noise, conditional copies are formally formulated in [11] as $\Omega=\Omega_{0}\times\widetilde{\Omega}_{1}$, where the original probability space follows the structure $\Omega=\Omega_{0}\times\Omega_{1}$, and $\widetilde{\Omega}_{1}$ represents the independent copy of the probability space $\Omega_{1}$. [11, Theorem 4.14, Volume II]
In the next section, we will provide the construction of conditional independent copies for the general case of a semimartingale $\bm{X}$ and an arbitrary sub $\sigma$-algebra ${\mathcal{G}}$ which is not necessarily associated with any probability space. As emphasized earlier, this construction is necessary for establishing Itô’s formula with the flow of conditional measures on semimartigales.
3.3. Construction of conditional independent copies: general cases
Given a probability space $(\Omega,{\mathcal{F}},{\mathbb{P}})$, let us consider the extended probability space $\overline{\Omega}=\Omega^{n+1}$ defined by

### v3

@@ 2.1. Assumptions and preliminary tools
For $p\geq 2$, a function $\Phi(\mu,\bm{y})$ is $C^{2,2}({\mathcal{P}}_{p}({\mathbb{R}}^{d})\times{\mathbb{R}}^{d})$ if for any $\mu\in\mathcal{P}_{p}({\mathbb{R}}^{d})$ and $\bm{y}\in{\mathbb{R}}^{d}$, there exists a continuous mapping $(\mu,\bm{y},\bm{x}_{1})\to\frac{\delta\Phi}{\delta\mu}(\mu,\bm{y},\bm{x}_{1})$, as well as a continuous mapping $(\mu,\bm{y},\bm{x}_{1},\bm{x}_{2})\to\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}(\mu,\bm{y},\bm{x}_{1},\bm{x}_{2})$ that is symmetric in its two arguments $(\bm{x}_{1},\bm{x}_{2})$ and with the following properties:
  • (Continuous differentiability) $\nabla_{\bm{y}}\Phi(\mu,\bm{y})$, $\nabla_{\bm{y}}^{2}\Phi(\mu,\bm{y})$, $\frac{\delta\Phi}{\delta\mu}(\mu,\bm{y},\bm{x}_{1})$, $\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}(\mu,\bm{y},\bm{x}_{1})$,
    $\nabla_{\bm{y}}\frac{\delta\Phi}{\delta\mu}(\mu,\bm{y},\bm{x}_{1})$ $\nabla^{2}_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}(\mu,\bm{y},\bm{x}_{1})$, $\nabla_{\bm{x}_{1}}\nabla_{\bm{y}}\frac{\delta\Phi}{\delta\mu}(\mu,\bm{y},\bm{x}_{1})$, $\nabla_{\bm{x}_{1}}\nabla_{\bm{x}_{2}}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}(\mu,\bm{y},\bm{x}_{1},\bm{x}_{2})$ all exist and are continuous for all $\bm{y},\bm{x}_{1},\bm{x}_{2}\in{\mathbb{R}}^{d}$, $\mu\in{\mathcal{P}}_{p}({\mathbb{R}}^{d})$.
@@
  • (Uniform polynomial-growth) there exists a constant $c>0$ such that for all $\bm{x}_{1},\bm{x}_{2},\bm{y}\in{\mathbb{R}}^{d}$, $\mu\in{\mathcal{P}}({\mathbb{R}}^{d})$,
@@
    $\bigg{|}\nabla^{2}_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\bigg{|}+\bigg{|}\nabla_{\bm{x}_{1}\bm{y}}\frac{\delta\Phi}{\delta\mu}\bigg{|}$ $\leq c\big{(}1+|\bm{x}_{1}|^{p-2}+|\bm{y}|^{{p-2}}\big{)},$
    $\bigg{|}\nabla_{\bm{x}_{1},\bm{x}_{2}}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\bigg{|}$ $\leq c\big{(}1+|\bm{x}_{1}|^{p-2}+|\bm{x}_{2}|^{p-2}+|\bm{y}|^{p-2}\big{)}.$
  • (Fundamental theorem of calculus) for everything $\mu,\nu\in{\mathcal{P}}_{2}({\mathbb{R}}^{d})$,
@@
    $\Phi(\mu,\bm{y})$- $\Phi(\nu,\bm{y})-\int_{0}^{1}\int_{{\mathbb{R}}^{d}}\frac{\delta\Phi}{\delta\mu}\big{(}\nu,\bm{y},\bm{x}_{1}\big{)}(\mu-\nu)({\mathrm{d}}\bm{x}_{1})$
    $=\int_{0}^{1}\int_{0}^{t}\int_{{\mathbb{R}}^{d}}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}s\mu+(1-s)\nu,\bm{y},\bm{x}_{1},\bm{x}_{2}\big{)}(\mu-\nu)({\mathrm{d}}\bm{x}_{1})(\mu-\nu)({\mathrm{d}}\bm{x}_{2}){\mathrm{d}}s{\mathrm{d}}t.$
Here $\frac{\delta\Phi}{\delta\mu}$ and $\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}$ are respectively (a version of) the first order and the second order linear derivatives of $\Phi$.
Remark 2.10.
Note that the linear derivative is in the sense of [21], which is properly defined up to an additive constant. This form of linear derivative is shown to be appropriate for analyzing the flow of measures on semimartingales [21, 31].
@@ 2.2. Main result
[…]
@@
$\qquad\quad+\frac{1}{2}\nabla^{2}_{\bm{x}_{1}\bm{x}_{2}}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-},\bm{X}^{\prime\prime}_{s-}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime\prime}]^{c}_{s}+\nabla^{2}_{\bm{x}_{1}\bm{y}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{Y}]^{c}_{s}\bigg{)}$
@@
for all $t\in[0,T]$, where $\bm{X}^{\prime},\bm{X}^{\prime\prime}$ are two conditionally independent copies of $\bm{X}$ given the sub $\sigma$-algebra ${\mathcal{G}}_{T}$ defined in some enlarged probability space $(\overline{\Omega},\overline{{\mathcal{F}}},\overline{{\mathbb{P}}})$, and $\overline{{\mathbb{E}}}[\cdot|{\mathcal{F}}]$ is the conditional expectation given $\mathcal{F}$} in the enlarged probability space.
@@
To ensure that the enlarged space $(\overline{\Omega},\overline{{\mathcal{F}}},\overline{{\mathbb{F}}},\overline{{\mathbb{P}}})$ is properly defined such that $\mu_{t},\bm{X}_{t},\bm{Y}_{t}$ and $\mathcal{F}$} are naturally extended to this enlarged space and $\bm{X}^{\prime}$$\bm{X}^{\prime\prime}$ are well-defined, one needs the precise notion of conditionally independent copies of stochastic processes. In addition, one needs to show the existence of these conditionally independent copies in Theorem 2.11. These will be studied in the next section and specified in Theorem 3.1 and Corollary 3.2.
Remark 2.13.
Theorem 2.11 includes the following Itô’s formulas as special cases:
  • The Itô’s formula for mean-field games and control problems with common noise developed in [9, 14]: take $X_{t}$ as an Itô’s process driven by both idiosyncratic noise $W$ and common noise $W^{0}$, let the common filtration be the filtration generated by $W_{0}$.
  • The Itô’s formula for flows of measures associated with possibly discontinuous processes without common noise studied in [31, 53]: set $X_{t}$ as a general semimartingale and choose $\mathcal{G}$} as the trivial $\sigma$-algebra.
  • The Itô’s formula for partially observed diffusions as in [4]: let $X_{t}$ be an Itô’s process and letting $\mathcal{G}$} be the filtration generated by the observation process. For more details, see Section 6 for an extension of this problem to jump-diffusion processes.
3. Conditionally Independent Copy of Stochastic Process
3.1. Why conditionally independent copy?
One of the key components in Theorem 2.11 is the notion of conditionally independent copies of stochastic processes. We now illustrate this concept through several examples in the context of flow of conditional laws for semimartingales and its role in the derivation of Itô’s formula.
Let us first consider a real-valued stochastic process $X_{t}$ of the form
@@ 3.1. Why conditionally independent copy?
where $W,W^{0}$ are two independent real valued Brownian motions, ${\mathbb{G}}=\{{\mathcal{G}}_{t}\}_{t\in[0,T]}$ is the filtration generated by $W^{0}$, ${\mathcal{G}}={\mathcal{G}}_{T}=\sigma(W^{0}_{u},u\leq T)$, and $\sigma_{t}$ and $\sigma_{t}^{0}$ are appropriate functions to ensure the well-definedness of the above SDE. By the classical Itô’s formula, the dynamic of $g(X_{t})$ for a smooth function $g$ is given by
@@
The above equation follows essentially from the fact that for progressively measurable process $(\eta_{t})_{t\in[0,T]},$
${\mathbb{E}}\bigg{[}\int_{0}^{t}\eta_{s}{\mathrm{d}}W_{s}\Big{|}{\mathcal{G}}\bigg{]}=0,\qquad{\mathbb{E}}\bigg{[}\int_{0}^{t}\eta_{s}{\mathrm{d}}W^{0}_{s}\Big{|}{\mathcal{G}}\bigg{]}=\int_{0}^{t}{\mathbb{E}}[\eta_{s}|{\mathcal{G}}]{\mathrm{d}}W^{0}_{s},\text{ for all }t\in[0,T].$ (3.2)
(See [47, Lemma B.1] for more details).
From (3.1) one can calculate the quadratic variation of $Z_{t}$:
@@
where $\widetilde{\sigma}_{t}^{0},\widetilde{X}_{t}$ are respectively conditionally independent copies of $\sigma^{0}_{t}$ and $X_{t}$ given $\mathcal{G}$}, as formally introduced in [14]. Moreover, one can derive for a smooth function $f$ the Itô’s formula for $f(Z_{t})$, which depends on the precise expression of
$\int_{0}^{t}f^{\prime}(Z_{s}){\mathrm{d}}Z_{s}=$ $\int_{0}^{t}f^{\prime}(Z_{s})\bigg{(}{\mathbb{E}}\Big{[}b_{s}g^{\prime}(X_{s})+\frac{1}{2}(\sigma_{s}^{2}+(\sigma^{0}_{s})^{2})g^{\prime\prime}(X_{s})\Big{|}{\mathcal{G}}\Big{]}{\mathrm{d}}s+{\mathbb{E}}\big{[}\sigma^{0}_{t}g^{\prime}(X_{s})|{\mathcal{G}}\big{]}{\mathrm{d}}W^{0}_{s}\bigg{)}$
@@
However, there is an issue when one considers a general continuous semimartingale $X$ and a general sub $\sigma$-algebra $\mathcal{G}$}, where the Itô’s formula takes the form:
@@
Indeed, in order to characterize the dynamics of $Z_{t}={\mathbb{E}}[g(X_{t})|{\mathcal{G}}]$, one needs an appropriate expression for ${\mathrm{d}}Z_{t}$, ${\mathrm{d}}\langle Z\rangle_{t}$, and more importantly for
[…]
@@
${\mathbb{E}}\bigg{[}\int_{0}^{t}\eta_{s}{\mathrm{d}}X_{s}\Big{|}{\mathcal{G}}\bigg{]}=\int_{0}^{t}{\mathbb{E}}[\eta_{s}|{\mathcal{G}}]{\mathrm{d}}{\mathbb{E}}[X_{s}|{\mathcal{G}}].$ (3.5)
Unfortunately this guess does not hold in general. For instance, take $X_{t}$ from (3.1) with $b_{t}=0$. By (3.2), the left hand side of (3.5) becomes
@@
and (3.5) holds if and only if $\eta_{s}$ and $\sigma_{s}^{0}$ are conditionally uncorrelated, given $\mathcal{G}$}.
[…]
$\int_{0}^{t}f^{\prime}(Z_{s}){\mathrm{d}}Z_{s}$ $=\int_{0}^{t}f^{\prime}(Z_{s}){\mathrm{d}}{\mathbb{E}}\big{[}g(X_{s})|{\mathcal{G}}\big{]}=\int_{0}^{t}f^{\prime}(Z_{s}){\mathrm{d}}\widetilde{\mathbb{E}}\big{[}g(X^{(1)}_{s})\big{]}=\widetilde{\mathbb{E}}\bigg{[}\int_{0}^{t}f^{\prime}(Z_{s}){\mathrm{d}}g(X^{(1)}_{s})\bigg{]}$
$=\widetilde{\mathbb{E}}\bigg{[}\int_{0}^{t}f^{\prime}(Z_{s})\Big{(}g^{\prime}(X^{(1)}_{s}){\mathrm{d}}X^{(1)}_{s}+\frac{1}{2}g^{\prime\prime}(X^{(1)}_{s}){\mathrm{d}}\langle X^{(1)}\rangle_{s}\Big{)}\bigg{]},$
@@
$\int_{0}^{t}f^{\prime\prime}(Z_{s}){\mathrm{d}}\langle Z\rangle_{s}$ $=\int_{0}^{t}f^{\prime}(Z_{s}){\mathrm{d}}\langle{\mathbb{E}}\big{[}g(X_{\cdot})|{\mathcal{G}}\big{]}\rangle_{s}=\widetilde{\mathbb{E}}\bigg{[}\int_{0}^{t}f^{\prime}(Z_{s}){\mathrm{d}}\langle g(X^{(1)}_{\cdot}),g(X^{(2)}_{\cdot})\rangle_{s}\bigg{]}$
$=\widetilde{\mathbb{E}}\bigg{[}\int_{0}^{t}f^{\prime\prime}(Z_{s})g^{\prime}(X^{(1)}_{s})g^{\prime}(X^{(2)}_{s}){\mathrm{d}}\langle X^{(1)},X^{(2)}\rangle_{s}\bigg{]},$
where $\widetilde{\mathbb{E}}$ denotes the expectation on $X^{(1)}$ and $X^{(2)}$ while keeping the information that is $\mathcal{G}$}-measurable. With these expressions, one can analyze (3.4) for a general semimartingale $X$, as will be detailed in (1) and (2) of Theorem 3.5.
Next, we will provide the construction of conditionally independent copies for the general case of a semimartingale $\bm{X}$ and an arbitrary sub $\sigma$-algebra $\mathcal{G}$} which is not necessarily associated with any random variable. As emphasized earlier, this construction is necessary for establishing Itô’s formula with the flow of conditional measures on semimartigales.
3.2. Construction of conditionally independent copies of stochastic process
Given a probability space $(\Omega,{\mathcal{F}},{\mathbb{P}})$, consider the extended probability space $\overline{\Omega}=\Omega^{n+1}$ defined by